# Do Similar Entities have Similar Embeddings?

### v1

@@ 5 Results and Discussion
5.1 Different Notions of Similarity
Table 2 reports results w.r.t. rank-based metrics (MRR, Hits@K) and RBO@K for all the KGEMs and datasets considered in this work.
@@ 5.1 Different Notions of Similarity
A crucial observation from our study is that the behavior of KGEMs varies not only across datasets but also across different classes within these datasets. For example, in Table 3, we clearly see that on YAGO4-19K, TransE captures the notion of SpanishMunicipality quite well compared to other models, but does not do well for MusicPlaylist entities (highlighted in blue). This means that different models are better suited for capturing the nuances of specific classes. Jain et al. [14] demonstrated that semantic representation in embeddings is not consistent across all KG entities, but is restricted to a small subset of them. Our results point in the same direction and suggest that KGEMs are not equipped with consistent capabilities to learn similar embeddings for entities within the same class. This problem appears at two different levels: First, we noted a general misalignment in the representation of certain classes across various datasets, regardless of the KGEM used. Second, there is noticeable inconsistency within specific classes, where entities might align well with their graph-based proximity in one KGEM but not in another.
@@
To generalize this observation, we computed correlation matrices for RBO values across models and datasets, focusing on class-specific performance. Our analysis, averaging these matrices, gives us a broader view of how different KGEMs perceive similarity on a per-class basis. Fig. 4 shows the rank correlations between models when considering 2-hop subgraphs, averaged over all datasets and RBO values. Notably, RDF2Vec stands out as it seems to capture a distinct notion of similarity compared to other KGEMs. It demonstrates that RDF2Vec fares better on different classes than other KGEMs. We also observe a high degree of variation between models of the same family: TransE and TransD expose a different notion of similarity (0.64), as RESCAL and TuckER do (0.6). Strong rank correlations are observed between DistMult and BoxE (0.85), DistMult and TransE (0.84), and BoxE and TransE (0.84) (Fig. 4). This demonstrates that while some KGEMs exhibit a close conceptualization of entity similarity, this is not universally the case. Therefore, RQ1 can be answered as follows: proximity in the embedding space does not consistently align with the notion of entity similarity in the KG, since this property substantially differs between models and datasets. Consequently, careful consideration is necessary when using KGEs for drawing conclusions about similar entities (e.g. for recommending items to similar users).
@@
Table 2: Rank-based and RBO results on the 6 datasets used in our experiments. As RDF2Vec is not designed for link prediction, it is not evaluated w.r.t. rank-based metrics. H@3 and H@10 stand for Hits@3 and Hits@10, respectively. RBO is abbreviated as R, while the number directly following it denotes whether 1-hop subgraphs or 2-hop subgraphs are considered. Bold fonts indicate which KGEM performs best for a given configuration (dataset-metric). Underlined results denote the second-best performing model.
@@
5.2 On the Correlation between Rank-based Metrics and RBO
@@
From a coarse-grained viewpoint, we note that MRR generally does not correlate well with RBO values. Two specific cases are that MRR is heavily (and positively) correlated with RBO@10 (1-hop) and RBO@100 (2-hop), which means that MRR can be seen as a reliable proxy measure for those two metrics (and vice versa). However, it still remains to be explained why the correlation between MRR and RBO@10 (1-hop) is so high, while a substantial drop in correlation between MRR and RBO@100 (1-hop) is observed. Another result to be assessed is why MRR correlates more with RBO@K with lower $K$ values when considering 1-hop subgraph and higher $K$ values when considering 2-hop subgraphs. Therefore, RQ2 cannot be answered in a conclusive way: in some special cases, values for rank-based metrics are heavily correlated with RBO values. But, as a whole, this statement is rarely observed. In other words: a good performance for link prediction does not imply that the also fulfills the KGE entity similarity assumption to a large extent, i.e., for use cases which heavily rely on entity similarity, choosing a KGE model by its performance on link prediction (which is often observed in research papers) might be a suboptimal strategy.
5.3 Analyzing Predicate Importance
@@
In what follows, we examine if that particular behavior may stem from KGEMs actually weighting predicates differently. We consequently pick random classes, and analyze how the predicate distribution of the top-K neighbors (in the embedding space) for entities of this class differs between models.
@@
Results for two picked classes are reported in Table 4, where each predicate is weighted in accordance to its frequency in the set of common triples between the 2-hop subgraphs of top-K neighbors in the embedding space (for $K=3$ and $K=10$) and a given entity of class MusicComposition and MusicGroup (YAGO4-19K). A class-based analysis reveals that predicates are differently ranked, i.e. different KGEMs may rely on different sets of predicates. For example, for $K=10$ neighbors and entities of the class MusicComposition, either isPartOf or inLanguage appears as the most frequent predicate depending on the KGEM considered.
We additionally observe that even when the ranking is the same between KGEMs – especially since the predicate distribution can be severely skewed towards one or few predicates – the relative importance of predicates can still differ. For instance, although genre is consistently ranked as the most frequent predicate for MusicGroup entities, results for RDF2Vec suggest a more balanced predicate distribution and a lesser emphasis put on this single predicate (highlighted in blue in Table 4). It is the only model to put a signficant emphasis on award (i.e., considers two MusicGroups as similar if they have won the same (1-hop) or similar (2-hop) awards). In some cases, a few KGEMs seem to assign a lower relevance to some specific predicates. For instance, under BoxE the top-3 and top-10 neighbors of entities that are MusicCompositions contain the predicate inLanguage at a much lesser frequency in their subgraphs’ intersection (highlighted in brown in Table 4), compared to other models. Answering RQ3, this suggests that a thorough analysis of predicate distribution seems essential if we want to understand the contributing factors to the notion of similarity for particular KGEMs.
@@ 6 Conclusion and Outlook
This work delved into the intricate relationship between entity similarity in KGs and their respective embeddings in different KGEMs, questioning the wide-spread KGE entity similarity assumption. Contrary to this belief, we showed that the choice of KGEM significantly influences the notion of entity similarity encoded in the resulting vector space. This finding has profound implications for a variety of downstream tasks where accurate entity similarity is crucial, e.g. recommender systems and semantic searches.
A common practice seen in many papers is to first compute different KGEMs on the KG at hand and evaluate them by their performance in link prediction, and then picking the one with the best performance for a downstream task relying on entity similarty. A critical takeaway from our study is that this common practice might not yield the best results in terms of capturing true entity similarity. Instead, our results advocate for a more nuanced approach. Moreover, in scenarios where the similarity of specific classes of entities is of paramount importance, a per-class analysis becomes essential. Such an approach allows for a tailored selection of KGEMs that are more adept at capturing the nuances and semantics of particular classes, thereby ensuring a more accurate and meaningful representation of entity similarity.

### v2

@@ 5 Results and Discussion
5.1 Different Notions of Similarity (RQ1)
Table 2 reports results w.r.t. rank-based metrics (MRR, Hits@K) and RBO@K for all the KGEMs and datasets considered in this work. Similar to MRR and Hits@K, RBO is computed for all entities in a KG and the average is reported.
@@ 5.1 Different Notions of Similarity (RQ1)
A crucial observation from our study is that the behavior of KGEMs varies not only across datasets but also across different classes within these datasets. Table 3 illustrates this insight on classes that were selected to provide a more intuitive understanding of how different KGEMs are differently able to capture similarities between instances of the same class. In particular, we clearly see that on YAGO4-19K, TransE captures the notion of SpanishMunicipality quite well compared to other models, but does not do well for MusicPlaylist entities (highlighted in blue). This means that different models are better suited for capturing the nuances of specific classes. Jain et al. [14] demonstrated that semantic representation in embeddings is not consistent across all KG entities, but is restricted to a small subset of them. Our results point in the same direction and suggest that KGEMs are not equipped with consistent capabilities to learn similar embeddings for entities within the same class. This problem appears at two different levels: First, we noted a general misalignment in the representation of certain classes across various datasets, regardless of the KGEM used. Second, there is noticeable inconsistency within specific classes, where entities might align well with their graph-based proximity in one KGEM but not in another.
@@
To generalize this observation, we computed correlation matrices for RBO values across models and datasets, focusing on class-specific performance. Our analysis, averaging these matrices, gives us a broader view of how different KGEMs perceive similarity on a per-class basis. Fig. 4 shows the rank correlations between models when considering 2-hop subgraphs, averaged over all datasets and RBO values. Notably, RDF2Vec stands out as it seems to capture a distinct notion of similarity compared to other KGEMs (cf. [23]). It demonstrates that RDF2Vec fares better on different classes than other KGEMs. We also observe a high degree of variation between models of the same family: TransE and TransD expose a different notion of similarity (0.64), as RESCAL and TuckER do (0.6). Strong rank correlations are observed between DistMult and BoxE (0.85), DistMult and TransE (0.84), and BoxE and TransE (0.84) (Fig. 4). This demonstrates that while some KGEMs exhibit a close conceptualization of entity similarity, this is not universally the case. Therefore, RQ1 can be answered as follows: proximity in the embedding space does not consistently align with the notion of entity similarity in the KG, since this property substantially differs between models and datasets. Consequently, careful consideration is necessary when using KGEs for drawing conclusions about similar entities (e.g. for recommending items to similar users).
@@
Table 2: Rank-based and RBO results. As RDF2Vec is not designed for link prediction, it is not evaluated w.r.t. rank-based metrics. H@3 and H@10 stand for Hits@3 and Hits@10, respectively. RBO is abbreviated as R, while the number directly following it denotes whether 1-hop subgraphs or 2-hop subgraphs are considered. Bold fonts indicate which KGEM performs best for a given configuration (dataset-metric). Underlined results denote the second-best performing model.
@@
5.2 Correlation between Rank-based Metrics and RBO (RQ2)
@@
From a coarse-grained viewpoint, we note that MRR generally does not correlate well with RBO values. Two specific cases are that MRR is heavily (and positively) correlated with RBO@10 (1-hop) and RBO@100 (2-hop), which means that MRR can be seen as a reliable proxy measure for those two metrics (and vice versa). However, it still remains to be explained why the correlation between MRR and RBO@10 (1-hop) is so high, while a substantial drop in correlation between MRR and RBO@100 (1-hop) is observed. Another result to be assessed is why MRR correlates more with RBO@K with lower $K$ values when considering 1-hop subgraph and higher $K$ values when considering 2-hop subgraphs. Therefore, RQ2 cannot be answered in a conclusive way: while in the general case, the correlation between performance in grouping similar entities and link prediction performance is moderate at best, some metrics show a higher correlation. This has a severe practical implication: the commonly observed practice to pick a KGE model which is good at link prediction for a task heavily relying on entity similarity has to be considered a suboptimal strategy.
5.3 Analyzing Predicate Importance (RQ3)
@@
To examine this behavior further, we pick random classes and count the predicates in the common triples of the top-K neighbors (in the embedding space) for entities of this class differs between models.
@@
Results for two picked classes are reported in Table 4, where each predicate is weighted in accordance to its frequency in the set of common triples between the 2-hop subgraphs of top-K neighbors in the embedding space (for $K=3$ and $K=10$) and a given entity of class MusicComposition and MusicGroup (YAGO4-19K). A class-based analysis reveals that predicates are differently ranked, i.e. different KGEMs may rely on different sets of predicates. For example, for $K=10$ and entities of the class MusicComposition, either isPartOf or inLanguage appears as the most frequent predicate depending on the KGEM considered.
We additionally observe that even when the ranking is the same between KGEMs – especially since the predicate distribution can be severely skewed towards one or few predicates – the relative importance of predicates can still differ. For instance, although genre is consistently ranked as the most frequent predicate for MusicGroup entities, results for RDF2Vec suggest a more balanced predicate distribution and a lesser emphasis put on this single predicate (highlighted in blue in Table 4). It is the only model to put a signficant emphasis on award (i.e., considers two MusicGroups as similar if they have won the same (1-hop) or similar (2-hop) awards). In some cases, a few KGEMs seem to assign a lower relevance to specific predicates. For instance, under BoxE the top-3 and top-10 neighbors of entities that are MusicCompositions contain the predicate inLanguage at a much lesser frequency in their subgraphs’ intersection (highlighted in brown in Table 4), compared to other models. Answering RQ3, we observe that the different notions of entity similarity reflected by different KGEMs are comparable, as most of them put a focus on the same set of predicates when determining entity similarity.
@@ 6 Conclusion and Outlook
This work delved into the intricate relationship between entity similarity in KGs and their respective embeddings in KGEMs, questioning the widespread KGE entity similarity assumption. Contrary to this belief, we showed that the choice of KGEM significantly influences the notion of entity similarity encoded in the resulting vector space. This finding has profound implications for a variety of downstream tasks where accurate entity similarity is crucial, e.g. recommender systems and semantic searches. For instance, if the proximity of graph embeddings does not align with our proposed metric for measuring entity similarity, the tacit assumption of downstream systems exploiting embeddings for recommender systems does not hold and caution is needed when deploying such systems.
A common practice is to evaluate KGEMs by their performance in link prediction, then picking the one with the best performance for a downstream task relying on entity similarty. A critical takeaway from our study is that this practice might not yield the best results in terms of capturing true entity similarity. Instead, our results advocate for cautiousness. Moreover, in scenarios where the similarity of specific classes of entities is of paramount importance, a per-class analysis becomes essential. This approach allows for a tailored selection of KGEMs that are more adept at capturing the nuances and semantics of particular classes, thereby ensuring a more accurate and meaningful representation of entity similarity.